OpenStudy (anonymous):

Using the equations expressing degrees Fahrenheit in terms of degrees Celsius (vice versa). At what temperature are the two scales equal? Please explain!

OpenStudy (amistre64):

when do the number for the degrees match?

OpenStudy (anonymous):

Are you studying linear equations?

OpenStudy (anonymous):

find points in terms of (Celsius, Fahrenheit) for freezing and boiling

OpenStudy (anonymous):

freezing is 0 celsius and 32 fahrenheit, boiling is 100 celsius and 212 F

OpenStudy (amistre64):

32 212 0 100 ------- 32,112 = 112/32 = 56/16 = 14/4; 7/2 .. doesnt look right

OpenStudy (anonymous):

so make two points and then work from those points like any linear equation problem

OpenStudy (anonymous):

no, the points should be C, F not C, C and F,F

OpenStudy (amistre64):

.... lol, thats right

OpenStudy (anonymous):

(0,32) and (100, 212)

OpenStudy (anonymous):

change in y over change in x for slope, then find the y0-intercept (it is one of the points - consider C to be x and F to be y)

OpenStudy (anonymous):

It still doesn't make sense. What would be the temperature if the equations for F and C are equal?